### v2

# Word Embeddings Revisited: Do LLMs Offer Something New?

Abstract
Learning meaningful word embeddings is key to training a robust language model. The recent rise of Large Language Models (LLMs) has provided us with many new word/sentence/document embedding models. Although LLMs have shown remarkable advancement in various NLP tasks, it is still unclear whether the performance improvement is merely because of scale or whether underlying embeddings they produce significantly differ from classical encoding models like Sentence-BERT (SBERT) or Universal Sentence Encoder (USE). This paper systematically investigates this issue by comparing classical word embedding techniques against LLM-based word embeddings in terms of their latent vector semantics. Our results show that LLMs tend to cluster semantically related words more tightly than classical models. LLMs also yield higher average accuracy on the Bigger Analogy Test Set (BATS) over classical methods. Finally, some LLMs tend to produce word embeddings similar to SBERT, a relatively lighter classical model.
1 Introduction
With the introduction of Word2Vec Mikolov et al. (2013a) and GLoVe Pennington et al. (2014), word embedding techniques and their semantic analysis have gained significant popularity in the field of deep NLP research. Since then, the size and complexity of the embedding models have increased dramatically, especially with the introduction of transformer-based architectures like BERT Devlin et al. (2018), BART Lewis et al. (2019), RoBERTa Liu et al. (2019), followed by rigorously fine-tuned sentence encoders like LASER Artetxe and Schwenk (2019), Universal Sentence Encoder (USE) Cer et al. (2018), SentenceBERT (SBERT) Reimers and Gurevych (2019), etc., and finally with the emergence of Large Language Models (LLMs) like GPT Brown et al. (2020), PaLM Chowdhery et al. (2022), LLaMA Touvron et al. (2023), etc. While transformer-based models not only offer embeddings for words but also for longer sequences like sentences and documents, learning meaningful word embeddings still remains fundamental for proper contextualization of longer sequences and robust language modeling.
While LLMs have shown remarkable success across various well-known NLP tasks Bubeck et al. (2023); Dai et al. (2022); Du et al. (2022); Smith et al. (2022); Sarkar et al. (2023); Akter et al. (2023), it remains an open question if their enhanced performance can solely be attributed to their larger scale or if the embeddings they generate are fundamentally distinct from those created by traditional encoding models such as Sentence-BERT or Universal Sentence Encoders Mahajan et al. (2023); Sarkar et al. (2022). Since the introduction of the Transformer architecture, model development has exploded with innovations, improving year after year. However, the medium has remained the same: written language. In theory, the similarity and relatedness of words have not changed, but the way that models treat words and their similarity has. More recently, Fournier et al. (2020) evaluated smaller transformer models like GPT-2 and BERT on analogy tasks and found that these models performed poorly on Encyclopedic and Lexicographic analogies, further raising the question of how embedding performance changes as model size increases. As such, we perform the following two analyses in this paper.
  • Word-Pair Similarity Analysis: Given two distinct word embedding techniques (classical or LLM-based), we compare and analyze their cosine similarity distributions of pairs of 1) Semantically Related, 2) Morphologically Related, and 3) Unrelated words, where the distributions are derived from their word embedding vectors.
  • Word Analogy Task Analysis: For a given word analogy task of the form $a:b::c:d$ (“a is to b as c is to d”), we compare the accuracies of various embedding techniques using vector arithmetic in the latent vector space.
2 Experiments and Results
2.1 Embeddings
A sampling of models from the last five years was used to create a multitude of embeddings for experimentation. We classify these models into two broad categories: Large Language Model (LLM) based – models with over 1B parameters and “Classical” models – those with under 1B parameters. Tested here are LLaMA2-7B (dim =4096=4096) from Meta AI Touvron et al. (2023), OpenAI’s embedding model ADA-002 (dim =1536=1536), and Google’s PaLM2 embedding model Gecko-001 (dim =768=768Anil et al. (2023). To more clearly see the differences between these models and older (“classical”) ones, Meta AI’s LASER (dim =1024=1024Artetxe and Schwenk (2019), Universal Sentence Encoder (USE) (dim =512=512Cer et al. (2018), and SentenceBERT (SBERT) (dim =384=384Reimers and Gurevych (2019) were also tested.
2.2 Dataset
In an extension of the original analogy task set out by Mikolov et al. (2013b), Gladkova et al. (2016) introduced Bigger Analogy Test Set (BATS), a larger set of related word pairs. This set is more comprehensive regarding the types of relations included and more forgiving in analogies with more than one correct answer Gladkova et al. (2016). This dataset forms the basis for the evaluations performed in this paper.
2.3 Analysis 1: Word-Pair Similarity
RQ-1: How do LLM-based embeddings differ from classical ones in terms of the expected cosine similarity for a randomly chosen pair of words?
To answer this question, we sampled a corpus of approximately $80,000$ distinct words from WordNet Fellbaum (1998) using all lemmas from all synsets. Next, we computed the cosine similarity of each pair of words using their vector representations extracted from each embedding model. The distribution of cosine similarities of all pairs of words ($\approx 6.4$ billion pairs) is shown in Figure 1. As seen, the mean and distribution for each model are meaningfully different, demonstrating significant differences in terms of their latent semantic spacing.
Figure 1: The distribution of cosine similarities between all pairs of words for each model.
Finding-1: LLM-based embeddings, particularly ADA and LLaMA, yield higher expected cosine similarity for a random pair of words than the same for PaLM and all classical embeddings.
RQ-2: Do LLM-based embeddings capture semantic similarity better than classical embeddings?
To answer this question, we used the BATS dataset, which provides pairs of related words across many categories describing how those words are related. These categories were used to create two distinct sets of word pairs: Morphologically Related Pairs and Semantically Related Pairs. We also include a third set, Uncategorized Pairs, which is made up of random word pairs. Next, we plot the distribution of cosine similarities between these pairs of words (by three categories) for all six embeddings (see Figure 2).
Figure 2 demonstrates that ADA, PaLM, and SBERT showed the biggest separation between related pairs (both morphologically and semantically related) and uncategorized pairs, which is a desired property. In all models, morphologically related words were more similar to each other than semantically related ones (somewhat expected). USE, LLaMA, and LASER struggled to distinguish between semantically related pairs and randomly chosen pairs.
Finding-2: LLMs are not always better than classical models in capturing semantic similarity (e.g., SBERT vs. LLaMA). Among LLMs, ADA and PaLM can effectively distinguish semantically related pairs from unrelated pairs, and so can SBERT. All models can effectively distinguish morphologically related pairs from unrelated pairs.
Figure 2: Histogram showing the distribution of cosine similarities for random, morphologically related, and semantically related pairs of words for each model.
2.4 Analysis 2: Word Analogy Task
RQ-3: Do LLMs improve the performance of word embeddings on word analogy tasks?
To answer this question, we followed the original word analogy task format set out by Mikolov et al. (2013b) and comprehensively evaluated the six embedding models on the word pairs from the BATS dataset. For words $a,b,c,d$, analogy $a:b::c:d$ and embedding function $f(x)$, it is expected that $f(b)-f(a)+f(c)\approx f(d)$, which we will refer to as the 3CosAdd method. Other approaches have been introduced for this task, including Pair Distance and 3CosMul (introduced by Levy and Goldberg (2014)). Later, Drozd et al. (2016) introduced new methods called 3CosAvg and LRCos, which achieved excellent performance in their experiments on classical models.
Method 3CosAdd 3CosAvg 3CosMul LRCos PairD SBERT 0.243 0.261 0.267 0.487 0.086 USE 0.174 0.212 0.187 0.450 0.025 LASER 0.227 0.260 0.237 0.284 0.121 ADA-002 0.412 0.447 0.424 0.375 0.232 LLaMA2 0.145 0.200 0.145 0.131 0.053 PaLM 2 0.398 0.458 0.417 0.534 0.193
Table 1: Performance on BATS Analogy. Blue denotes the best accuracy; bold black denotes the second best.
For all methods, the 3 words used as the input for the analogy were excluded from the answers, and top-1 accuracy was measured. For fairness, the same corpus from Section 2.3 was used for each model, and the arithmetic results for each method were used to find the nearest neighbor in the corpus. These results are shown in Table 1, with the best-performing embedding for each method shown in blue. Both ADA and PaLM performed very well, while LLaMA had the worst performance in the LLM category. Among classical embeddings, SBERT performed quite well, often ranked as the third best. Full information about each model’s accuracy in each category can be found in the appendix.
Figure 3: For each model, the cosine similarity of related words was found and ranked according to all pairs of words. Here, the difference in ranking between model pairs for certain BATS categories is shown.
Finding-3: Among LLMs, ADA and PALM perform significantly better than classical models on word analogy tasks, while SBERT (a classic model) is often ranked as third, suggesting that SBERT can be an efficient alternative in resource-constrained devices.
(a) Kendal $\tau$
(b) Spearman $\rho$
Figure 4: Correlation coefficients for each pair of models, found using a large dataset of pairs of words.
2.5 Do LLMs Offer Something New?
In order to determine whether LLMs offer something new/very different, we investigated how embedding models (LLMs and Classical) agreed or disagreed with each other regarding the similarity of word pairs. To be more specific, we computed the difference in rank between two related words for each pair of embedding models. For example, if the 5th closest word to “bad” according to ADA-002’s embedding was “worst”, while "worst" was the 10th nearest word to “bad” according to LLaMA’s embeddings, we would compute a difference of 5-5 for that word pair while comparing ADA vs. LLaMA. If two models tend to agree on the relative similarity of word pairs most of the time, we would expect a mean of 0 for this value and a small variance.
Figure 3 shows the results using the BATS corpus, which allows us to conclude that all models except LLaMA agree with each other reasonably well on the similarity of words related by morphology. However, for semantically related words, we see much more disagreement, but we see that two LLMs, ADA and PaLM, tend to agree with each other and share some agreement with SBERT. This is surprising, as ADA and PaLM are two of the heaviest models, while SBERT is one of the lightest. We found very little agreement between other model pairs for other categories. All models except LASER tended to rank the related words higher than LLaMA in nearly all categories. LASER and LLaMA disagreed very strongly across all categories, suggesting that LASER and LLaMA embeddings are quite different from all others.
For a more robust evaluation of agreement among models, we computed statistical measures of correlation. First, all the cosine similarities of all pairs of words from the Section 2.3 corpus were computed for each embedding model. Then, the correlation between two different embedding models was computed in terms of word pair similarities. Table 4 shows both Kendall’s $\tau$ and Pearson’s $\rho$ between each pair of embeddings. These results show that, interestingly, SBERT and ADA-002 are the most semantically similar models, while USE and ADA are the most different.
Finding-4: Two of the LLMs, PaLM and ADA, tended to agree with each other, but they also surprisingly meaningfully agreed with SBERT, again suggesting that SBERT can be an efficient alternative when resources are constrained.
3 Conclusion
In this paper, we investigated the compelling question of whether LLM-based word embeddings significantly differ from classical embedding techniques like SBERT and USE. Using WordNet and the popular BATS dataset, we created a corpus of 80,000 unique words and 6.4 billion word pairs. We analyzed six popular embedding models (three LLM-based and three classical models) in terms of two properties: 1) the similarity of their latent semantic spaces, and 2) their performance on word analogy tasks. Our results suggest that PaLM and ADA, two LLM-based models, tend to agree with each other and yield the highest performance on word analogy tasks, but they also surprisingly meaningfully agree with SBERT. This suggests that while PaLM and ADA can capture meaningful semantics and yield high accuracy, SBERT can be an efficient alternative when resources are constrained.
4 Limitations
This work only analyzes six models due to time and resource constraints. More work will need to be done to increase the number of models tested to draw a more general conclusion on the difference between Classical and LLM-based embeddings.
Numerous works Linzen (2016); Fournier et al. (2020) have illuminated issues in using the standard analogy task to determine if semantic information is encoded in word embeddings. Therefore, this paper has refrained from making statements implying that one embedding is ‘better’ than another.
This work also relies upon the assumption that cosine similarity is a reasonable metric to compute semantic similarity between two word embedding vectors. While the interpretation of cosine similarity can itself be debatable, cosine similarity is still the most popular similarity metric used in NLP literature. However, our findings are limited by the inherent limitations of the cosine similarity metric.
Appendix A Appendix
See Table 2 for a more granular description of the performance of each model on specific categories of BATS.
Model Analogy Method
1. Inflectional
Morphology
2. Derivational
Morphology
3. Encyclopedic
Semantics
4. Lexicographic
Semantics
LLAMA2-7B 3CosAdd 0.230 0.271 0.055 0.023
LLAMA2-7B 3CosAvg 0.326 0.362 0.086 0.026
LLAMA2-7B 3CosMul 0.230 0.276 0.053 0.022
LLAMA2-7B LRCos 0.150 0.148 0.176 0.050
LLAMA2-7B PairDistance 0.066 0.130 0.013 0.001
LLAMA2-7B SimilarToAny 0.065 0.043 0.037 0.011
LLAMA2-7B SimilarToB 0.130 0.118 0.054 0.026
PaLM-Gecko-001 3CosAdd 0.743 0.609 0.118 0.122
PaLM-Gecko-001 3CosAvg 0.794 0.668 0.232 0.136
PaLM-Gecko-001 3CosMul 0.768 0.648 0.128 0.124
PaLM-Gecko-001 LRCos 0.780 0.714 0.404 0.238
PaLM-Gecko-001 PairDistance 0.466 0.249 0.048 0.008
PaLM-Gecko-001 SimilarToAny 0.165 0.027 0.011 0.035
PaLM-Gecko-001 SimilarToB 0.270 0.082 0.030 0.108
GPT-ADA-002 3CosAdd 0.761 0.677 0.115 0.097
GPT-ADA-002 3CosAvg 0.802 0.734 0.148 0.102
GPT-ADA-002 3CosMul 0.776 0.697 0.122 0.100
GPT-ADA-002 LRCos 0.606 0.482 0.280 0.132
GPT-ADA-002 PairDistance 0.546 0.323 0.052 0.006
GPT-ADA-002 SimilarToAny 0.155 0.044 0.005 0.029
GPT-ADA-002 SimilarToB 0.276 0.134 0.038 0.090
LASER 3CosAdd 0.431 0.434 0.022 0.022
LASER 3CosAvg 0.484 0.506 0.030 0.020
LASER 3CosMul 0.448 0.454 0.023 0.023
LASER LRCos 0.510 0.482 0.116 0.028
LASER PairDistance 0.230 0.245 0.009 0.003
LASER SimilarToAny 0.087 0.027 0.004 0.007
LASER SimilarToB 0.198 0.072 0.012 0.020
USE 3CosAdd 0.397 0.156 0.039 0.103
USE 3CosAvg 0.442 0.190 0.084 0.132
USE 3CosMul 0.436 0.165 0.049 0.100
USE LRCos 0.722 0.412 0.396 0.270
USE PairDistance 0.076 0.012 0.008 0.005
USE SimilarToAny 0.101 0.032 0.006 0.035
USE SimilarToB 0.204 0.098 0.026 0.098
SBERT 3CosAdd 0.461 0.393 0.046 0.073
SBERT 3CosAvg 0.474 0.418 0.058 0.092
SBERT 3CosMul 0.506 0.424 0.062 0.074
SBERT LRCos 0.808 0.642 0.270 0.228
SBERT PairDistance 0.135 0.184 0.021 0.003
SBERT SimilarToAny 0.178 0.065 0.003 0.019
SBERT SimilarToB 0.302 0.154 0.020 0.088
Table 2: BATS performance across categories with methods.